# EEGDM: Learning EEG Representation with Latent Diffusion Model

### v2

@@ III-C2 Conditional Guidance
III-C3 Conditional Input Setup
@@
The treatment of conditional inputs and the initialization process significantly influence the generation quality in diffusion models. We adopt the adaLN-Zero block from DiT, which has shown superior performance, to handle the conditional inputs (the EEG representations $e$ and the noise timestep $t$). Specifically, we regress the scale and shift parameters $\gamma$ and $\beta$ in the Transformer blocks based on the sum of the $e$ and $t$. Additionally, we regress the dimension-wise scaling parameters $\alpha$, which are applied just before each residual connection in the DiT blocks. The MLP for all $\alpha$ is initialized to output the zero vector, which effectively initializes the entire DiT block as an identity function. In addition to layer normalization, we incorporate conditional embeddings into residual connections, which enhance the vanilla adaLN and improve the quality of the generated EEG signals.
IV Experiment
IV-A Datasets
@@
Pre-training datasets. We utilize the Temple University Hospital EEG Corpus (TUEG) [obeid2016temple] to pre-train the proposed EEGDM. TUEG is a very large publicly available clinical EEG dataset containing clinical-grade recordings collected in real-world hospital environments. It has over 40 different channel configurations and varying temporal durations of recordings. The majority of the recordings are acquired at a sampling rate of 256 Hz. TUEG exhibits substantial heterogeneity in patient demographics, channel configurations, and recording equipment, making it particularly suitable for pre-training our model.
Datasets of downstream specific tasks. (1) TUEV (event type classification) [harati2015improved]: An EEG dataset containing six event categories: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG). The EEG signals are recorded at 23 channels and 256 Hz sampling rate. (2) TUAB (abnormal detection) [lopez2015automated]: A corpus of EEG recordings annotated as either normal or abnormal. Similar to TUEV, the EEG signals of TUAB are recorded at 23 channels and 256 Hz sampling rate. (3) SHU-MI (motor imagery classification) [ma2022large]: An EEG motor imagery dataset designed for decoding imagined movements. It features 32 EEG channels sampled at 250 Hz, including two types of motor imagery tasks: left hand and right hand. (4) CL-Drive (cognitive load assessment) [angkan2024multimodal]: It is a multimodal dataset designed for cognitive load assessment, which includes EEG signals. The EEG data are recorded using four channels at a sampling frequency of 256 Hz. This study explores the ternary classification of cognitive load into high, medium, and low levels. (5) Seed-VII (emotion recognition) [jiang2024seed]: It is a large-scale multimodal emotion dataset. Seed-VII includes 62-channel EEG and eye-movement data from 20 healthy subjects across seven emotion categories (disgust, fear, sad, neutral, happy, anger, surprise). The preprocessed EEG signals were downsampled to 200 Hz. (6) DMER (emotion distribution prediction) [yang2024multimodal]: It is a multimodal emotion recognition corpus that includes EEG signals along with other modalities (GSR, PPG, and frontal facial videos). In each trial, participants watched video clips and subsequently completed a self-assessment using a ten-item version of the Positive Affect and Negative Affect Schedule (PANAS) [mackinnon1999short], reporting the intensity of ten emotional adjectives. Each emotion rating ranges from 1 (not at all) to 5 (extremely), and the resulting scores are transformed into a 10-dimensional emotion distribution. The EEG recordings comprise 18 channels sampled at 300 Hz.
@@
It is worth noting that we primarily select downstream datasets from the past five years, a strategic shift from the benchmarks used in works like CBraMod [wang2024cbramod] or EEGPT [wangeegpt]. This decision stems from the need to evaluate the model against the fidelity and complexity of modern neuroscience standards. Older datasets often lack the high spatial density and signal quality required to stress-test a foundation model, whereas recent paradigms — ranging from naturalistic stimuli to multimodal retrieval — better assess the capacity for generalizable, high-level cognitive decoding. However, to ensure fair comparison with established baselines, we make exceptions for the TUAB and TUEV datasets. Despite falling outside the five-year window, these two widely adopted benchmarks are retained as standardized anchors, allowing us to validate our model’s performance on classical clinical tasks alongside its capabilities on modern, complex scenarios.
IV-B Experimental Setup
@@
For the pre-training dataset TUEG, following prior work [wang2024cbramod], we select 19 common EEG channels. All EEG signals are resampled to 200 Hz and segmented into non-overlapping 30-second samples (19 channels × 6,000 time points). Each EEG patch is set to a temporal length of 1 second, resulting in 570 patches per sample. For all downstream datasets, we maintain consistency with our pre-training data by resampling EEG signals to 200 Hz and setting each EEG patch duration to 1 second.
For the downstream datasets TUEV and TUAB, consistent with previous studies [wangeegpt, jiang2024large, wang2024cbramod], we utilize 16 common bipolar montage channels from the international 10-20 system, with sample durations of 5 seconds and 10 seconds, respectively. We adopt the train-test splits provided by these datasets. For the CL-Drive dataset, we use the channel configuration and sample duration (10 seconds) provided by the dataset and apply 10-fold subject-wise cross-validation. For the SHU-MI, DMER, and SEED-VII datasets, we use the channel configurations provided by each dataset and segment the signals into 4-second samples. We employ a subject-independent split strategy for these datasets. We fine-tune EEGPT [wangeegpt], LaBraM-Base [jiang2024large] (the only variant with available pre-trained weights), and CBraMod [wang2024cbramod] based on their publicly available code and pre-trained weights, except when their experimental results are already reported in the original papers.
For the EEG generator (diffusion module) and encoder, we adopt the DiT and ViT architectures, respectively. Following common practices in generative modeling literature, we maintain an exponential moving average (EMA) of model weights with a decay rate of 0.9999 during pre-training. The specific hyperparameters for pre-training are detailed in Table I.
We evaluate the representation learning capabilities of EEGDM through fine-tuning on downstream tasks. Specifically, we predict corresponding labels using a linear predictor based on the outputs of the EEG encoder. Following the common evaluation protocols, we present results under the fine-tuning setting, where the predictor is jointly trained with the EEG encoder (i.e., the pre-trained encoder is not frozen). We employ cross-entropy loss for downstream classification tasks and KL divergence loss for downstream distribution prediction tasks. The hyperparameters for fine-tuning EEGDM on downstream tasks are provided in Table II. We conduct a series of quantitative and qualitative experiments to evaluate the proposed model, demonstrating its strong representation capabilities across diverse downstream datasets.
TABLE I: Hyperparameters for EEGDM pre-training.
Hyperparameter Value
Optimization Batch size 32
Peak learning rate 5e-4
Minimal learning rate 5e-5
Learning rate scheduler OneCycleLR
Weight decay 0.05
EMA decay rate 0.9999
Gradient clipping norm 1
Optimizer AdamW
${\beta}_{1}$ 0.9
${\beta}_{2}$ 0.999
EEG Generator Layers N 8
Hidden size 512
Heads 8
Normalization type LayerNorm
EEG Encoder Layers 8
Hidden size 512
MLP size 2048
Heads 8
Sampling Classifier-free guidance 9
Diffusion training steps 1000
TABLE II: Hyperparameters for downstream fine-tuning. ($\star$) Depends on the dataset.
Hyperparameter Value
Batch size 64
Peak learning rate 5e-4
Minimal learning rate 5e-5
Learning rate scheduler OneCycleLR
Optimizer AdamW
${\beta}_{1}$ 0.9
${\beta}_{2}$ 0.999
Weight decay 0.05
Total epochs 30, 150
Warmup Epochs 6, 30
IV-C Results and Analysis
In this section, we show that EEGDM achieves competitive performance across diverse downstream tasks, demonstrating the effectiveness of diffusion-based representation learning for EEG analysis. All experiments are conducted with multiple random seeds, and we report the mean and standard deviation of evaluation metrics. The reported model size corresponds to the downstream networks used for fine-tuning, as the generative module is not involved in downstream tasks. Results are compared against baselines, as shown in Tables III, IV, V and VI. Compared to other approaches, EEGDM consistently maintains either the best or second-best performance in key metrics (balanced accuracy or KL divergence).
@@ IV-C Results and Analysis
TABLE III: Performance comparison on TUEV. The best metric is highlighted in bold, and the second-best metric is underlined.
@@
[…]
TABLE IV: Performance comparison on TUAB. The best metric is highlighted in bold, and the second-best metric is underlined.
@@
[…]
TABLE V: Performance comparison on DMER. The best metric is highlighted in bold, and the second-best metric is underlined.
@@
TABLE VI: The results of different methods on various datasets. The best metric is highlighted in bold, and the second-best metric is underlined.
@@
TABLE VII: The results of the ablation studies for PCA and EEG augmentation, using the metric of average balanced accuracy.
@@
On the TUEV dataset, EEGDM achieves a balanced accuracy of 68.19%, outperforming both the previous state-of-the-art method CBraMod (66.71%) and other baselines. This superiority highlights the exceptional discriminative power of the representations learned by the diffusion model for classifying different event types in EEG signals.
Table IV shows that EEGDM achieves a balanced accuracy of 81.80% and an AUC-PR of 90.85% on the TUAB dataset. Securing the second-best performance on those metrics, EEGDM overall demonstrates comparable performance with a competitive margin. The consistent performance across both balanced accuracy and AUC-PR metrics suggests that EEGDM learns robust representations that generalize well for binary abnormal detection tasks, maintaining high discriminative power even in challenging clinical scenarios.
Table V presents the results for the DMER dataset, which uses distributional metrics. EEGDM achieves best performance across various divergence and similarity metrics, with the exception of the Clark metric, where it ranks second. The overall high performance on DMER, a task requiring the prediction of a continuous emotion distribution rather than a discrete class, strongly suggests that the diffusion-based pre-training learns rich and fine-grained features that preserve the complex, subtle, and continuous emotional states embedded in the EEG signals.
Table VI further validates the generalizability and robustness of EEGDM across diverse BCI tasks. Specifically, on the SHU-MI dataset, EEGDM ranks second in terms of balanced accuracy (63.24%) and AUC-PR (70.70%), while achieving the best performance in AUROC (69.96%), demonstrating competitive performance compared with CBraMod. For the CL-Drive dataset, EEGDM shows the best balanced accuracy and the best Cohen’s Kappa, demonstrating superior cognitive load feature extraction capabilities. On the Seed-VII dataset, EEGDM continues its strong performance with the best performance in balanced accuracy and ranks second in weighted F1 and Cohen’s Kappa, further validating its effectiveness in emotion-related tasks.
These comprehensive results across diverse datasets validate the effectiveness of using EEG signal generation as a self-supervised pre-training objective, demonstrating that the combination of encoder and diffusion denoiser, enhanced by PCA and channel augmentation, successfully learns high-quality EEG representations for improved downstream task performance.
@@
Figure 3: Performance comparison with different numbers of PCA components on TUEV and TUAB datasets.
IV-D Ablation Study
We conduct ablation studies to analyze the contributions of PCA and channel augmentation components, respectively, by setting up four variant versions as shown in Table VII, in which the average balanced accuracy on three random seeds is reported. Variant D consistently achieves the best performance across all settings, demonstrating that PCA and EEG augmentation are important for our model. However, we find that using PCA or EEG augmentation alone is either detrimental to or has minimal impact on our model’s performance. This may be attributed to the fact that PCA reduces the dimensionality of the source signal, making the model more sensitive to the diversity of conditional information, while data augmentation provides sufficient diversity to support learning in this reduced-dimensional space. Consequently, the PCA and EEG augmentation operations achieve complementary effects in terms of generalization performance and accuracy. Furthermore, for TUAB, we randomly selected 10% of the training set for fine-tuning and tested it on the entire test set (see the last column in Table VII); we also use this setting in Fig. 3. The results show that the advantages of incorporating PCA and EEG augmentation remain evident even when small-scale datasets are used for fine-tuning.
@@ IV-D Ablation Study
We also investigate the impact of the number of PCA components on downstream task performance. Fig. 3 presents results on TUEV and TUAB with varying numbers of PCA components, which reveal a nonlinear relationship between the number of PCA components and downstream task performance. In particular, insufficient number of components may result in large information loss, whereas an excessive number introduces noise contamination, both of which degrade model performance. To balance these trade-offs, we empirically select an intermediate value of 20 components, which yields optimal performance in our experiments.

### v3

@@ 3.3.2. Conditional Guidance
3.3.3. Conditional Input Setup
@@
The treatment of conditional inputs and the initialization process significantly influence the generation quality in diffusion models. We adopt the adaLN-Zero block from DiT, which has shown superior performance, to handle the conditional inputs (the EEG representations $e$ and the noise timestep $t$). Specifically, we regress the scale and shift parameters $\gamma$ and $\beta$ in the Transformer blocks based on the sum of the $e$ and $t$. Additionally, we regress the dimension-wise scaling parameters $\rho$, which are applied just before each residual connection in the DiT blocks. The MLP for all $\rho$ is initialized to output the zero vector, which effectively initializes the entire DiT block as an identity function. In addition to layer normalization, we incorporate conditional embeddings into residual connections, which enhance the vanilla adaLN and improve the quality of the generated EEG signals.
4. Experiment
4.1. Datasets
@@
Pre-training datasets. We utilize the Temple University Hospital EEG Corpus (TUEG) (Obeid and Picone, 2016) to pre-train the proposed EEGDM. TUEG is a large-scale publicly available clinical EEG dataset containing clinical-grade recordings collected in real-world hospital environments. It encompasses over 40 different channel configurations and recordings of varying temporal durations, with the majority acquired at a sampling rate of 256 Hz. TUEG exhibits substantial heterogeneity in patient demographics, channel configurations, and recording equipment, making it particularly suitable for pre-training our model.
Downstream tasks and Datasets. To comprehensively evaluate our method, we consider multiple downstream tasks, including event type classification (TUEV dataset (Harati et al., 2015)), abnormal detection (TUAB dataset (Lopez et al., 2015)), motor imagery classification (SHU-MI dataset (Ma et al., 2022)), cognitive load assessment (CL-Drive dataset (Angkan et al., 2024)), emotion recognition (Seed-VII dataset (Jiang et al., 2024b)), and emotion distribution prediction (DMER dataset (Yang et al., 2024b)). Detailed descriptions of these datasets are provided in the supplementary material.
@@
Notably, we primarily select downstream datasets from the past five years, representing a strategic shift from the benchmarks used in prior works such as CBraMod (Wang et al., 2025) and EEGPT (Wang et al., 2024). This choice is motivated by the need to evaluate the model against the fidelity and complexity of contemporary neuroscience standards. Older datasets often lack the high spatial density and signal quality required to stress-test a foundation model, whereas recent paradigms—ranging from naturalistic stimuli to multimodal retrieval—offer a more effective assessment of the model’s capacity for generalizable, high-level cognitive decoding. Nevertheless, to ensure fair comparison with established baselines, we make exceptions for the TUAB and TUEV datasets. Although these two widely adopted benchmarks fall outside the five-year window, they are retained as standardized anchors, allowing us to validate our model’s performance on classical clinical tasks alongside its capabilities on modern, complex scenarios.
4.2. Experimental Setup
@@
Data Pre-processing. For the pre-training dataset TUEG, following prior work (Wang et al., 2025), we select 19 common EEG channels. All EEG signals are resampled to 200 Hz and segmented into non-overlapping 30-second samples (19 channels $\times$ 6,000 time points). Each EEG patch is set to a temporal length of 1 second, resulting in 570 patches per sample. For all downstream datasets, we maintain consistency with our pre-training setup by resampling EEG signals to 200 Hz and setting each EEG patch duration to 1 second.
@@
Table 1. Performance comparison on TUEV. The best result is highlighted in bold, and the second-best result is underlined.
@@
Table 2. Performance comparison on TUAB. The best result is highlighted in bold, and the second-best result is underlined. NA indicates that the result is not available in the original literature (Wang et al., 2024).
@@
Baselines and Implementation. We compare our EEGDM against state-of-the-art self-supervised EEG representation learning frameworks, including EEGPT (Wang et al., 2024), LaBraM-Base (Jiang et al., 2024c) (the only variant with available pre-trained weights), CBraMod (Wang et al., 2025), and Puah-EEGDM-SSM (Puah et al., 2025). All baselines are fine-tuned using their publicly available code and pre-trained weights, except in cases where results were directly reported in the original papers. For our model, we adopt the DiT architecture for the EEG generator (diffusion module) and a ViT backbone for the encoder. The model is pre-trained on a cluster of four NVIDIA A100 GPUs with a batch size of 32. The entire pre-training phase encompasses 329,400 steps, totaling approximately 112.96 wall-clock hours.
Evaluation Protocols. We evaluate our method under two settings: fine-tuning and linear probing (with a frozen encoder), to comprehensively assess the quality of the learned representations. Extensive experiments are conducted to evaluate the proposed model, demonstrating its strong representation capability across diverse downstream datasets.
Detailed hyperparameter configurations, data splitting strategies, and implementation specifics for both pre-training and downstream evaluation protocols are provided in the supplementary material.
Table 3. Performance comparison on DMER. The best result is highlighted in bold, and the second-best result is underlined.
@@
[…]
Table 4. Performance comparison on SHU-MI, CL-Drive and Seed-VII datasets. The best result is highlighted in bold, and the second-best result is underlined. Results for Puah-EEGDM-SSM on SHU-MI are omitted due to performance near chance levels.
@@
[…]
4.3. Results and Analysis
In this section, we demonstrate that EEGDM achieves competitive performance across diverse downstream tasks, underscoring the effectiveness of diffusion-based representation learning for EEG analysis. All experiments are conducted with multiple random seeds, and we report the mean and standard deviation of evaluation metrics. A comprehensive evaluation reveals that among all compared methods, our EEGDM is the only model that attains a top-two position on all six downstream datasets (except for EEGDM (T-Pre) on TUEV) with respect to the primary evaluation metrics (balanced accuracy and KL divergence).
As shown in Table 1, EEGDM achieves a balanced accuracy of 68.19% on the TUEV dataset when pre-trained on the large-scale TUEG corpus, outperforming several state-of-the-art baselines. Notably, while the general-purpose version of EEGDM trails the results reported by Puah-EEGDM-SSM (Puah et al., 2025) (75.57%) on this specific benchmark, it is essential to consider the difference in pre-training data distributions. Unlike our model, which targets universal feature extraction across diverse domains (as evidenced by its superior results across other downstream datasets), Puah-EEGDM-SSM was pre-trained and evaluated exclusively on the TUEV dataset, benefiting from a high degree of domain alignment. To ensure a rigorous and fair architectural comparison, we introduced a task-specific variant of our model, denoted as EEGDM (T-Pre), which follows the same protocol as Puah-EEGDM-SSM (Puah et al., 2025) by utilizing TUEV for both pre-training and downstream evaluation. Under this aligned setting, EEGDM (T-Pre) achieves a state-of-the-art accuracy of 79.71%. This result demonstrates the superior representational capacity of our Latent Diffusion Transformer (DiT) architecture operating on a refined manifold. It further indicates that the performance gap in the general setting is a byproduct of domain shift in pre-training data rather than architectural limitations. The discrepancy between balanced accuracy and F1-score suggests that EEGDM’s latent space prioritizes holistic representation over majority-class optimization, leading to better pattern diversity at the cost of higher sensitivity to class boundary noise.
Table 2 shows that EEGDM achieves a balanced accuracy of 81.80% and an AUC-PR of 90.85% on the TUAB dataset. Securing the second-best performance on those metrics, EEGDM demonstrates competitive performance with a narrow margin. The consistent performance across both balanced accuracy and AUC-PR metrics suggests that EEGDM learns robust representations that generalize well to binary abnormality detection tasks, maintaining high discriminative power even in challenging clinical scenarios.
Table 3 presents the results for the DMER dataset, which uses distributional metrics. EEGDM achieves the best performance across various divergence and similarity metrics, with the exception of the Clark metric, where it ranks second. The strong performance on DMER, a task requiring the prediction of an emotion distribution rather than a discrete class, strongly suggests that diffusion-based pre-training learns rich, fine-grained features that preserve the complex, subtle, and continuous emotional states embedded in EEG signals.
Table 4 further validates the generalizability and robustness of EEGDM across diverse BCI tasks. Specifically, on the SHU-MI dataset, EEGDM ranks second in balanced accuracy (63.24%) and AUC-PR (70.70%), while achieving the best performance in AUROC (69.96%), demonstrating competitive performance compared with CBraMod. Notably, the results for Puah-EEGDM-SSM (Puah et al., 2025) on this specific benchmark are omitted from the table, as its performance fluctuated near chance levels (e.g., accuracy $\approx 50\%$). This discrepancy suggests that while domain-specific pre-training (as used in (Puah et al., 2025)) may excel in aligned tasks, our Latent Diffusion Transformer (DiT) architecture, pre-trained on a broader corpus, exhibits superior cross-task adaptability, particularly in capturing the subtle spatial-temporal patterns required for motor imagery. On the CL-Drive dataset, EEGDM achieves the best balanced accuracy and the best Cohen’s Kappa, demonstrating superior cognitive load feature extraction capabilities. On the Seed-VII dataset, EEGDM continues its strong performance with the best balanced accuracy and ranks second in weighted F1 and Cohen’s Kappa, further validating its effectiveness in emotion-related tasks.
Table 5. Performance comparison of two data augmentation strategies on the CL-Drive dataset.
Strategy Balanced Accuracy ↑ Weighted F1 ↑ Cohen’s Kappa ↑
[…]
Beyond representation learning, a distinctive advantage of our proposed EEGDM is its generative capability to synthesize high-fidelity EEG data, an ability inherently absent in mask-reconstruction frameworks such as MAE. This ability is particularly effective in mitigating the pervasive issues of data sparsity and class imbalance in neural decoding. We evaluate this capability on the CL-Drive dataset—a relatively small-scale benchmark characterized by significant class skewness—using two augmentation strategies: isoportion expansion and balanced augmentation. As summarized in Table 5, our generative augmentation significantly improves downstream performance. By populating under-represented manifolds with synthetic yet semantically consistent samples, EEGDM effectively rectifies biased decision boundaries caused by data scarcity. These results underscore that our diffusion-based pre-training not only distills robust features but also serves as a powerful plug-and-play data engine, enhancing the model’s adaptability to real-world, long-tailed EEG distributions.
Comprehensive qualitative evaluations, including t-SNE visualization and EEG signal reconstruction, are provided in the supplementary material to further validate the representation quality and generative fidelity of EEGDM.
4.4. Ablation Study
To verify the individual and collective contributions of PCA and channel augmentation, we evaluate four variants of EEGDM, as summarized in Table 6. Our analysis reveals a compelling synergistic relationship between these two components that extends beyond simple additive effects. Interestingly, the results indicate that employing either PCA or channel augmentation in isolation provides marginal improvements or even leads to slight performance degradation. While PCA projects signals into a latent space with improved signal-to-noise ratio (SNR), the reduction in dimensionality may inadvertently discard fine-grained temporal nuances. Without sufficient input diversity, the encoder tends to underfit the complex underlying brain states. Conversely, although channel augmentation enriches the training perspectives, it significantly increases the optimization burden when the model is tasked with reconstructing raw, noisy signals.
The state-of-the-art performance achieved by Variant D underscores a powerful complementary effect. PCA simplifies the latent manifold, making the model more sensitive to the structural information provided by the augmented views. Meanwhile, the diverse augmented inputs prevent the model from losing critical information—effectively compensating for the potential information loss incurred by dimensionality reduction. To further assess the generalizability of this mechanism, we conducted a low-resource fine-tuning experiment on the TUAB dataset using only 10% of the training data (last column); we also use this setting in Fig. 3. The results consistently show that the integration of PCA and EEG augmentation remains vital even in small-scale data settings.
@@ 4.4. Ablation Study
Table 6. The results of the ablation studies for PCA and EEG augmentation, using the metric of average balanced accuracy.
@@
Figure 3. Performance comparison with different numbers of PCA components on TUEV and TUAB datasets.
@@
We also investigate the impact of the number of PCA components on downstream task performance. Fig. 3 presents results on TUEV and TUAB with varying numbers of PCA components, revealing a non-monotonic relationship between the number of PCA components and downstream task performance. Specifically, an insufficient number of components may result in substantial information loss, whereas an excessive number introduces noise contamination, both of which degrade model performance. To balance these trade-offs, we empirically select an intermediate value of 20 components, which yields optimal performance in our experiments.
In summary, our experimental results highlight three key advantages of EEGDM: (1) High-quality EEG generation, as validated by its effectiveness in augmenting an imbalanced and small-scale dataset; (2) Effective representation learning, as demonstrated by competitive linear probing results and substantial performance gains through fine-tuning; and (3) Competitive performance across diverse downstream tasks, as evidenced by consistent improvements in balanced accuracy and robustness over state-of-the-art baselines.
5. Conclusion
We present EEGDM, a novel latent diffusion framework for EEG representation learning, through explicit conditioning on EEG channel augmentation. Our approach adopts the generative task as the self-supervised EEG representation learning objective. By (1) integrating a denoiser with the EEG encoder and (2) leveraging PCA-based projection and EEG channel augmentation, EEGDM performs the generation task in the latent space using rich conditional information while simultaneously learning powerful EEG semantic representations that can be applied to diverse downstream tasks. Comprehensive experimental results demonstrate the model’s effectiveness across multiple downstream tasks.
Supplementary Material
Appendix A Dataset Details for Downstream Tasks
This section provides a detailed description of the downstream datasets employed to evaluate the performance of our proposed model.
(1) TUEV (event type classification) (Harati et al., 2015): An EEG dataset containing six event categories: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG). The EEG signals are recorded with 23 channels at a sampling rate of 256 Hz.
(2) TUAB (abnormal detection) (Lopez et al., 2015): A corpus of EEG recordings annotated as either normal or abnormal. Similar to TUEV, the EEG signals in TUAB are recorded with 23 channels at a sampling rate of 256 Hz.
(3) SHU-MI (motor imagery classification) (Ma et al., 2022): An EEG motor imagery dataset designed for decoding imagined movements. It features 32 EEG channels sampled at 250 Hz, including two types of motor imagery tasks: left hand and right hand.
(4) CL-Drive (cognitive load assessment) (Angkan et al., 2024): A multimodal dataset designed for cognitive load assessment, which includes EEG signals. The EEG data are recorded using four channels at a sampling frequency of 256 Hz. This study explores the ternary classification of cognitive load into high, medium, and low levels.
(5) Seed-VII (emotion recognition) (Jiang et al., 2024b): A large-scale multimodal emotion dataset. Seed-VII includes 62-channel EEG and eye-movement data from 20 healthy subjects across seven emotion categories (disgust, fear, sad, neutral, happy, anger, surprise). The preprocessed EEG signals were downsampled to 200 Hz.
(6) DMER (emotion distribution prediction) (Yang et al., 2024b): A multimodal emotion recognition corpus that includes EEG signals along with other modalities (GSR, PPG, and frontal facial videos). In each trial, participants watched video clips and subsequently completed a self-assessment using a ten-item version of the Positive Affect and Negative Affect Schedule (PANAS) (Mackinnon et al., 1999), reporting the intensity of ten emotional adjectives. Each emotion rating ranges from 1 (not at all) to 5 (extremely), and the resulting scores are transformed into a 10-dimensional emotion distribution. The EEG recordings comprise 18 channels sampled at 300 Hz.
Appendix B Detailed Experimental Settings for Downstream Datasets
For the downstream datasets TUEV and TUAB, consistent with previous studies (Wang et al., 2024; Jiang et al., 2024c; Wang et al., 2025), we utilize 16 common bipolar montage channels from the international 10-20 system, with sample durations of 5 seconds and 10 seconds, respectively. We adopt the train-test splits provided by these datasets. For the CL-Drive dataset, we use the channel configuration and sample duration (10 seconds) provided by the dataset and apply 10-fold cross-validation. For the SHU-MI, DMER, and SEED-VII datasets, we use the channel configurations provided by each dataset and segment the signals into 4-second samples. A subject-independent split strategy is adopted for these datasets.
Appendix C Implementation Details
Table 7. Hyperparameters for EEGDM pre-training.
Hyperparameter Value
Optimization Batch size 32
Peak learning rate 5e-4
Minimal learning rate 5e-5
Learning rate scheduler OneCycleLR
Weight decay 0.05
EMA decay rate 0.9999
Gradient clipping norm 1
Optimizer AdamW
$\beta$} (0.9, 0.999)
Diffusion Module Layers N 8
Hidden size 512
Heads 8
Normalization type LayerNorm
EEG Encoder Layers 8
Hidden size 512
MLP size 2048
Heads 8
Sampling Classifier-free guidance 9
Diffusion training steps 1000
Pre-training: The pre-training of EEGDM is conducted on the large-scale TUEG dataset to learn robust, general-purpose neural representations. We utilize the AdamW optimizer with a peak learning rate of 5e-4, which is decayed to a minimal value of 5e-5 following a OneCycleLR scheduling policy. We apply an Exponential Moving Average (EMA) on the model weights with a decay rate of 0.9999. The architecture consists of an 8-layer EEG Encoder and an 8-layer Diffusion Module, both featuring a hidden size of 512 and 8 attention heads. Detailed hyperparameter configurations for the optimization process, model architecture, and diffusion sampling are summarized in Table 7.
Downstream tasks: We evaluate the pre-trained EEGDM across diverse downstream scenarios using two protocols: fine-tuning and linear probing. In the fine-tuning setting, both the pre-trained encoder and the task-specific prediction head are jointly optimized. For linear probing, the EEG encoder remains strictly frozen, and only the prediction head is trained to assess the quality of the learned representations. To accommodate the varying scales and complexities of different downstream datasets, we employ task-specific hyperparameter configurations, as detailed in Table 8. All downstream experiments are conducted using the AdamW optimizer to ensure training stability. Detailed hyperparameters for both protocols, including batch sizes, learning rates, and weight decay, are summarized in Table 8.
Table 8. Hyperparameters for downstream fine-tuning and linear probing. ($\star$) Depends on the dataset.
Hyperparameter Fine-tuning Linear Probing
Encoder status Unfrozen Frozen
Learning rate 5e-4 2e-4, 5e-4, 2e-3
Total epochs 30, 150 30, 50, 100
Weight decay 0.05 0.05, 0.02, 0.01
Batch size 64 32, 64
Optimizer AdamW AdamW
Learning rate scheduler OneCycleLR OneCycleLR, ReduceLROnPlateau, CosineAnnealingLR
$\beta$} (0.9, 0.999) (0.9, 0.999)
Warmup epochs 6, 30 3, 10
Appendix D Qualitative Analysis
To evaluate the quality of the learned representations, we visualize the EEG data embeddings using t-SNE (Van der Maaten and Hinton, 2008) as presented in Fig. 4. We randomly sample EEG signals from the test sets of two downstream tasks in a class-balanced manner and extract their representations using our pre-trained encoder. The t-SNE visualizations exhibit well-separated clusters across all datasets, indicating that our EEGDM learns discriminative representations capable of capturing task-relevant patterns. This clear clustering structure further demonstrates the EEGDM’s ability to learn generic yet informative EEG representations that can be successfully adapted to diverse downstream tasks.
To further validate the generative quality, we provide visualizations of reconstructed signals and their Power Spectral Density (PSD). As shown in Fig. 5, the reconstructed EEG signals on the pre-training dataset, as well as their frequency-domain representations, closely match the ground truth. While minor discrepancies remain at the fine-grained level, the global temporal structure and frequency characteristics are well preserved. These results indicate that the proposed model can generate realistic EEG signals and effectively capture high-level semantic information from EEG data.